# NAP-Tuning: Neural Augmented Prompt Tuning for Adversarially Robust Vision-Language Models

### v1

@@ V-F Perturbation Magnitude
TABLE III: Accuracy (%) under $\ell_{\infty}$-bounded adversarial perturbations. Models are trained and evaluated with varying perturbation magnitudes (measured in $\epsilon/255$). Bold values indicate the highest robustness at each test-time perturbation level.
@@
V-G Few-shot Learning
We examine the impact of shot count on model performance. Fig. 6 illustrates the relationship between number of shots (1-16) and model accuracy across diverse datasets. Clean accuracy shows consistent improvement, with average accuracy increasing from 40.6% (single shot) to 73.2% (16 shots)—a 1.8× improvement. More remarkably, robust accuracy exhibits steeper relative gains, rising from 14.1% to 49.8%—a dramatic 3.5× improvement.
This differential impact reveals a key insight: while few-shot learning is known to enhance clean accuracy, its benefits for robust performance are substantially more pronounced. The results indicate that robustness requires a richer representation of class concepts that becomes increasingly available with additional examples. These findings align with our observations in Section V-C, confirming that increasing shot count offers tremendous potential for performance enhancement in our method, providing a straightforward path to significantly improved robustness without architectural modifications.
@@ V-G Few-shot Learning
V-H Context Vector Count
We examine how the number of learnable context vectors ($\mathcal{V}_{t}$) affects model performance. Fig. 7 shows results for models with 1-16 context vectors on ImageNet and Oxford Flowers datasets. For ImageNet, clean accuracy remains relatively stable (55.9%-58.1%) across different vector counts, while robust accuracy shows modest improvements with increased vectors, peaking at 29.6% with 14-16 vectors. On Oxford Flowers, both metrics maintain high performance across all configurations, with clean accuracy ranging from 95.5% to 96.5% and robust accuracy between 85.2% and 86.7%. These findings suggest that the number of context vectors $\mathcal{V}_{t}$ has a relatively minor impact on overall performance compared to the influence of layer count in Multi-Layer Prompts observed in Section V-C. This indicates that input-level representational capacity is not the primary performance bottleneck, particularly for less complex datasets where even few context vectors prove sufficient.
@@ VI Conclusion and Discussion
This paper presents the Neural Augmentor framework for Multi-modal Adversarial Prompt Tuning (NAP-Tuning), a novel architecture-focused approach to enhancing adversarial robustness in Vision-Language Models. Our work addresses significant limitations in existing adversarial prompt tuning methods by fundamentally reconceptualizing prompt architecture through feature purification mechanisms.
VI-A Conclusion
We have extended adversarial prompt tuning from text-only to multi-modal contexts and from single-layer to multi-layer architectures while introducing token refiners that enable modality-specific and layer-specific feature correction. This architectural innovation represents a paradigm shift from previous loss-focused approaches, emphasizing the critical role of feature purification in addressing adversarial distortions. Through extensive experimentation, we have demonstrated that NAP-Tuning substantially outperforms existing methods, achieving state-of-the-art robustness against strong attacks while maintaining competitive clean accuracy.
VI-B Broader Implications
Our findings have several significant implications: 1) highlight the insufficiency of merely transplanting prompt tuning techniques from generalization domains to adversarial defense without architectural reconsideration; 2) underscore the importance of addressing adversarial perturbations at the feature level rather than solely through loss modification.
VI-C Discussion
Future work could explore adaptive token refiners that dynamically adjust their behavior based on detected perturbation characteristics, integration of uncertainty quantification mechanisms to improve refinement reliability, and extension to other multi-modal tasks beyond classification. Additionally, investigating the theoretical connections between feature purification and information bottleneck principles could yield deeper insights into adversarial robustness in representation learning.
In conclusion, our work establishes that addressing adversarial robustness in VLMs requires fundamental architectural innovations rather than merely adapting existing prompt tuning techniques. The Neural Augmentor framework provides a principled approach to this challenge, offering both improved performance and conceptual insights that advance our understanding of robust multi-modal learning.

### v2

@@ V-D1 Analysis of Feature Purification
TABLE VI: Accuracy (%) under $\ell_{\infty}$-bounded adversarial perturbations. Models are trained and evaluated with varying perturbation magnitudes (measured in $\epsilon/255$).
@@
Refer to caption Refer to caption Refer to caption Refer to caption
Original Adversarial AdvPT AdvPT-V
Refer to caption Refer to caption Refer to caption Refer to caption
AdvPT-VLI APD FAP NAP-Tuning
Figure 9: EigenCAM visualization of attention patterns. NAP-Tuning maintains focus on semantically meaningful regions (cat’s face) even under adversarial attack, while other methods exhibit significant attention shifts toward irrelevant areas.
Table V quantifies the feature distortion magnitude under attack. A distinct capability hierarchy emerges from the results. Shallow approaches (Vanilla CLIP, AdvPT) exhibit significant distortion, confirming that input-level optimization alone fails to restrain feature corruption. Methods incorporating deeper interventions (e.g., APD, AdvPT-V) achieve moderate mitigation. In sharp contrast, NAP-Tuning outperforms all competitors by a wide margin, achieving the lowest distortion score. This represents a drastic reduction compared to the baseline, validating our method’s unique ability to maintain structural integrity in the embedding space.
These quantitative findings are strongly corroborated by the EigenCAM [19] visualizations in Fig. 9. While baseline models suffer from severe attention dispersion—scattering focus across irrelevant background regions—NAP-Tuning maintains precise semantic consistency. Specifically, our model retains focus on discriminative regions (e.g., facial features) even under adversarial attack, closely mirroring the attention patterns of clean inputs. This alignment between reduced feature distortion and preserved attention focus confirms our central hypothesis: adversarial vulnerability stems primarily from internal feature corruption, which the Neural Augmentor effectively rectifies.
V-D2 Analysis of Model-Image Defense Integration
We investigate NAP-Tuning’s capability to integrate with image-based defense mechanisms. Similar to AdvPT [37], our approach focuses exclusively on model-side adaptation, allowing it to complement image-based defense techniques. This presents an advantage over methods that already incorporate image transformations in their design, which often cannot be effectively combined with other preprocessing defenses.
Figure 10: Performance comparison across eleven datasets demonstrates consistent improvements when combining model-centric NAP-Tuning with the image-processing DiffPure method.
Figure 11: Effect of adversarial regularization weight ($\alpha$) on clean and robust accuracy on Flowers dataset.
To demonstrate this complementary capability, we evaluate NAP-Tuning combined with DiffPure [20], a diffusion-based adversarial purification technique. Fig. 10 shows the combined approach consistently outperforms standalone NAP-Tuning, with an average improvement of 12.9 percentage points. These results highlight a key advantage of our model-focused approach: NAP-Tuning enhances network parameter robustness while remaining fully compatible with input-processing techniques. This orthogonality enables effective integration that addresses different aspects of adversarial vulnerability. The consistent improvements across diverse datasets suggest that combining complementary defense strategies represents a promising direction for building more robust visual recognition systems.
V-E Parameter Sensitivity Analysis
Finally, we evaluate the sensitivity of NAP-Tuning to key hyperparameters and training settings. These experiments validate the robustness and stability of our approach across different configurations.
V-E1 Impact of Adversarial Regularization ($\alpha$)
We investigate the sensitivity of NAP-Tuning to the adversarial regularization weight ($\alpha$) in Fig. 11. The analysis reveals a highly favorable decoupling of metrics: clean accuracy remains remarkably stable across the entire parameter sweep, showing negligible fluctuation regardless of the regularization strength. In contrast, robust accuracy exhibits a rapid ascent as $\alpha$ increases from minimal values, effectively stabilizing around $\alpha=2.0$. Beyond this threshold ($\alpha\geq 5.0$), performance saturates at a high level with diminishing returns. Consequently, we adopt $\alpha=5.0$ as our default configuration. This setting yields a near-optimal balance without strictly necessitating the absolute peak observed at higher values (e.g., $\alpha=20.0$). These findings confirm that our approach is insensitive to precise hyperparameter tuning, maintaining consistent effectiveness across a broad operational range.
@@ V-E1 Impact of Adversarial Regularization ($\alpha$)
V-E2 Impact of Perturbation Magnitude ($\epsilon$)
We evaluate the interplay between perturbation budgets and model performance in Table VI. A clear trade-off emerges from the data: increasing the training perturbation magnitude ($\epsilon$) substantially bolsters robustness against stronger attacks, yet this hardening incurs a discernible cost in clean accuracy. Specifically, models trained with larger budgets demonstrate vastly superior resilience when subjected to matching or higher-intensity threats. In contrast, models trained on weaker perturbations suffer from rapid degradation when facing attacks that exceed their training budget. This aligns with well-established observations in traditional adversarial training, confirming that robust prompt tuning follows the same fundamental principle: resistance to extreme perturbations is achieved by sacrificing a degree of standard generalization.
V-E3 Impact of Few-shot Count
We examine the impact of shot count on model performance. Fig. 12 illustrates the relationship between number of shots (1-16) and model accuracy across diverse datasets. Clean accuracy shows consistent improvement, with average accuracy increasing from 40.6% (single shot) to 73.5% (16 shots). More remarkably, robust accuracy exhibits steeper relative gains, rising from 14.1% to 49.9%. This differential impact reveals a key insight: while few-shot learning is known to enhance clean accuracy, its benefits for robust performance are substantially more pronounced. The results indicate that robustness requires a richer representation of class concepts that becomes increasingly available with additional examples. These findings align with our observations in Section V-C1, confirming that increasing shot count offers tremendous potential for performance enhancement in our method, providing a straightforward path to significantly improved robustness without architectural modifications.
V-E4 Impact of Context Vector Count
We investigate the sensitivity of model performance to the length of learnable context vectors ($\mathcal{V}_{t}$) in Fig. 13. On the complex ImageNet dataset, clean accuracy remains remarkably stable regardless of vector length, while robust accuracy exhibits only marginal gains as the number of vectors increases, eventually saturating at higher counts. This stability is even more pronounced on the simpler Flowers dataset, where both metrics maintain consistently high performance with negligible variance across all configurations. Crucially, these fluctuations are minimal when compared to the significant impact of prompt depth observed in Section V-C1. This confirms that input-level representational capacity is not the primary bottleneck; rather, the structural depth of the intervention plays a far more decisive role in determining robustness than the mere quantity of input tokens.
@@ VI Conclusion and Discussion
This paper presents the Neural Augmentor framework (NAP-Tuning) as a modular structural augmentation for enhancing the adversarial robustness of Vision-Language Models. By introducing internal TokenRefiners into a coordinated multi-modal and multi-layer prompting structure, we enable the explicit rectification of adversarial distortions within the feature manifold. Our empirical results across 11 benchmarks demonstrate that this internal purification mechanism significantly outperforms conventional input-side alignment methods, particularly under rigorous evaluation protocols. Crucially, the plug-and-play nature of the Neural Augmentor ensures high compatibility with frozen VLM backbones, providing a scalable and parameter-efficient defense for robust multi-modal learning.
Our analysis underscores that adversarial vulnerability in VLMs stems primarily from internal feature distortions that input-side prompt tuning alone cannot fully rectify. By shifting the defense focus to structural feature-level intervention, NAP-Tuning provides a principled way to restore representational integrity. Future research will focus on developing adaptive TokenRefiners capable of dynamic adjustment based on perturbation intensity, as well as extending this purification logic to generative multimodal architectures beyond contrastive learning paradigms.